Circular No. 7639

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2001ai, 2001aj, AND 2001ak A. V. Filippenko and R. Chornock, University of California at Berkeley,

report that CCD spectra (range 390–1000 nm), obtained on May 30 UT with the Keck-II 10-m reflector, reveal that SN 2001aj (IAUC 7606) is a type-II supernova beginning to enter the nebular phase (NED recession velocity 7940 km/s), SN 2001ak (IAUC 7607) is a type-II supernova in the nebular phase (NED recession velocity 5280 km/s), and SN 2001ai (IAUC 7605) is a type-Ic supernova ∼ 2 months past maximum brightness (in agreement with the report of Matheson et al., IAUC 7605).

SUPERNOVA 2001bg IN NGC 2608 S. Benetti, G. Altavilla, E. Cappellaro, A. Pastorello, and M. Turatto,

Padova; R. Cosentino and A. Zacchei, Telescopio Nazionale Galileo (TNG); I. J. Danziger, Trieste; and F. Patat, European Southern Observatory, communicate: “On May 12.92 UT, we obtained a high-resolution spectrum (range 464–782 nm, resolution 0.011 nm) with the TNG (+ SARG) of the type-Ia SN 2001bg (cf. IAUC 7626). The spectrum shows Na I D absorptions originating both in the Milky Way (equivalent width, EW = 0.026 nm for the doublet) and in two extragalactic clouds with recession velocities of 2136 and 2171 km/s, respectively — likely associated with the supernova’s parent galaxy (the recession velocity for NGC 2068 archived in NED is 2135 km/s). The total Na I D EW from the extragalactic clouds is 0.125 nm. Schlegel et al. (1998, Ap.J. 500, 525) give E(B − V) = 0.039 for the galactic absorption in the direction of NGC 2068. Assuming that the (Na I D)-dust ratio for the extragalactic clouds is the same as for the galactic component, we estimate E(B − V) ∼ 0.19 for the extinction of SN 2001bg in the parent galaxy. The presence of possible nearby dust clouds makes SN 2001bg a good candidate for the search of a late light echo. We found no sign in the spectrum of narrow Hα absorptions/emissions (EW < 0.007 nm for unresolved features).”

CCD R photometry by K. Hornoch, Lelekovice, Czech Republic: May 10.836 UT, 13.78; 11.919, 13.74; 13.876, 13.77; 20.847, 14.01; 25.843, 14.27.

V1548 AQUILAE = NOVA AQUILAE 2001 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V1548 Aql has been given to this nova (cf. IAUC 7627, 7628).

2001 May 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green